OpenStudy (anonymous):

Evaluate the function for the given value of X f(x)=x+3 when x=-3

OpenStudy (anonymous):

Everytime you get a question like this, substitute every x for the number you got to evaluate, in this case -3. so we say f(-3) = (just put -3 where you have a x) and you get :\[f(-3) = -3+3 = 0\]

OpenStudy (anonymous):

thanks :) but what if it was f(x)=x squared -3x when x=3

OpenStudy (anonymous):

3^2=9 f(9)=9+3=12

OpenStudy (anonymous):

no, you will get :\[3^2 -3\times3 = 9 - 9 = 0\]

OpenStudy (anonymous):

-3 times 3 is -9 because a negative number times a positive number is negative. Remember opposite signs means the result is negative. Same signs means its positive.

OpenStudy (anonymous):

oh my bad I thought the same obvious equation

OpenStudy (anonymous):

okay lets do one together f(x)= |x-2| when x=5 whats the best way to start?

OpenStudy (anonymous):

I5-2I=I3I=3